Problem 12-16 (Part Level Submission) The following information (in $ millions) was taken from the recent annual reports of tProblem 12-17 (Part Level Submission) Poston, Inc., and Victor Enterprises are competitors in the toy industry. Last year theSelected account information in $ millions) from Target Corporations 2015 annual report follows. While the company has authoPippin Piping Co. Comparative Income Statements For the Years Ended December 31 2016 2015 Sales revenue $576,600 $523,000 Cos2015 $25,800 127,000 80,700 10,000 243,500 576,000 $819,500 2014 $20,500 94,300 56,000 4,600 175,400 540,300 $715,700 PippinPrepare common-size financial statements for the three years. (Round percentage answers to 1 decimal places, e.g. 32.7. ListAssets 2016 2015 2014 Liabilities & Stockholders EquityStockholders Equity
Answer #1
Company A Company B
Cost of goods sold=(A) $ 27,303.10 $     26,558.00
Average Inventory=(B) $    2,758.10 $       4,949.50
Inventory Turnover=Cost of goods sold/Average Inventory
Inventory Turnover=(A)/(B) 9.9 Times 5.4 Times
Target Corporation
2015 2014
Income from Continuing operation=(A) $    3,321.00 $   2,449.00
Total Assets $ 40,262.00 $ 41,172.00
Total common stockholder's equity $ 12,957.00 $ 13,997.00
Average total Assets=(Beginning Assets+Ending Assets)/2
Average total Assets=($41172+$40262)/2=(B) $ 40,717.00
Return on Total Assets=Income from Continuing operation/Average total assets 8.2%
Return on Stockholder's Equity=Income from continuing operation/Average Stockholder's equity
Income from Continuing operation=(A) $    3,321.00
Average Stockholder's Equity=(Beginning Stockholder's equity+Ending Stockholder's Equity)/2=($13997+$12957)/2=(B) $ 13,477.00
Return on Stockholder's Equity=(A)/(B) 24.6%